### v1

# Some addition theorems for spin-weighted spherical harmonics

Abstract
We present some addition theorems for spin-weighted spherical harmonics, generalizing previous results for scalar (spin-zero) spherical harmonics. These addition theorems involve sums over the azimuthal quantum number of products of two spin-weighted spherical harmonics at different points on the two-sphere, either (or both) of which are differentiated with respect to one of their arguments.
I Introduction
Spin-weighted spherical harmonics (SWSH) [1, 2] are a class of functions on the two-sphere, generalizing the standard (spin-00) spherical harmonics [3]. While originally introduced in the context of studies of the BMS group [1, 2], SWSH have found application in a wide range of fields [4], including the study of anisotropies in the cosmic microwave background [5, 6, 7, 8, 9, 10]; gravitational physics, including gravitational waves and perturbations of black hole space-times [11, 12, 13, 14]; geosciences [15, 16]; anisotropic turbulence [17]; and electromagnetism [18]. Further details of the properties of SWSH and a more extensive list of references can be found in, for example, [15, 16], see also [19, 20, 21, 22, 23, 24] for more details.
In this paper we present some generalizations of the well-known addition theorem for SWSH [5, 12, 15, 16], which we have been unable to find explicitly in the literature to date. It should be emphasized that our focus in this paper is SWSH functions, rather than scalar, tensor or spinor harmonics or their generalizations, for which some addition theorems can be found in [25, 26].
The outline of this work is as follows. Sec. II reviews the key definitions and properties of SWSH which are required for our analysis. For ease of reference, our new addition theorems are presented in Sec. III, prior to their derivation in Sec. IV. Our conclusions can be found in Sec. V.
II Properties of SWSH
The SWSH ${}_{s}Y_{\ell}^{m}(\theta,\phi)$ are functions on the two-sphere ${\mathbb{S}}^{2}$ which have spin-weight $s$, where $s$ is an integer or half-integer [1, 2]. We employ the usual spherical polar coordinates $(\theta,\phi)$ on ${\mathbb{S}}^{2}$, with $\theta\in[0,\pi]$ and $\phi\in[0,2\pi)$. The polar angle $\theta$ is the angle between the vector joining the point on the surface of ${\mathbb{S}}^{2}$ to the origin and the $z$-axis; while $\phi$ is the azimuthal angle in the $(x,y)$-plane, so that a point on ${\mathbb{S}}^{2}$ has Cartesian coordinates $(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$.
The SWSH depend on three quantum numbers $s$, $\ell$ and $m$. The spin $s$ is a positive or negative integer or half-integer. The orbital angular momentum quantum number $\ell$ then takes values $\ell=|s|,|s|+1,|s|+2,\ldots$, and the azimuthal quantum number $m$ takes the values $m=-\ell,-\ell+1,\ldots\ell-1,\ell$.
There are a number of different definitions of the SWSH in the literature. In this work, we follow the conventions of [15], in particular:
Definition 1.
The SWSH ${}_{s}Y_{\ell}^{m}(\theta,\phi)$ are defined in terms of the Wigner $D$-matrices by [15]
${}_{s}Y_{\ell}^{m}(\theta,\phi)=\left(-1\right)^{s}{\sqrt{\frac{2\ell+1}{4\pi}}}D^{\ell*}_{m,-s}(\phi,\theta,0),$ (1)
where * denotes complex conjugation.
Remark.
Our results in this paper are valid for both integer and half-integer spins $s$. When $s$ is a half-integer, the factor of $(-1)^{s}$ in (1) will be purely imaginary (see, for example, [24] for some specific examples of SWSH in this case).
Our derivation of new addition theorems rests on the action of the operators ð${}_{s}\eth\,$ and ð¯${}_{s}{\overline{\eth}}\,$, as defined below.
Definition 2.
The operators ð${}_{s}\eth\,$ and ð¯${}_{s}{\overline{\eth}}\,$ act on the SWSH as follows [1, 2, 15]:
ð${}_{s}\eth\,{}_{s}Y_{\ell}^{m}(\theta,\phi)$ $=-\left[\partial_{\theta}+\frac{{\rm{i}}}{\sin\theta}\partial_{\phi}-s\cot\theta\right]{}_{s}Y_{\ell}^{m}(\theta,\phi),$ (2a)
ð¯${}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}(\theta,\phi)$ $=-\left[\partial_{\theta}-\frac{{\rm{i}}}{\sin\theta}\partial_{\phi}+s\cot\theta\right]{}_{s}Y_{\ell}^{m}(\theta,\phi).$ (2b)
Remark.
To avoid confusion, we denote by ðssubscriptsuperscriptitalic-ð${}_{s^{\prime}}\eth\,^{\prime}$, ð¯ssubscriptsuperscript¯italic-ð${}_{s^{\prime}}{\overline{\eth}}\,^{\prime}$ the operators ð${}_{s}\eth\,$ and ð¯${}_{s}{\overline{\eth}}\,$ with $(\theta,\phi)$ replaced by $(\theta^{\prime},\phi^{\prime})$ and with spin $s^{\prime}$ instead of $s$.
Proposition 3.
The operators ð${}_{s}\eth\,$, ð¯${}_{s}{\overline{\eth}}\,$ respectively increase and decrease the spin-weight $s$ of an SWSH ${}_{s}Y_{\ell}^{m}$ by one [1, 2, 15, 16]:
ð${}_{s}\eth\,{}_{s}Y_{\ell}^{m}$ $={}_{s+1}Y_{\ell}^{m}{\sqrt{(\ell-s)(\ell+s+1)}},$ (3a)
ð¯${}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}$ $=-{}_{s-1}Y_{\ell}^{m}{\sqrt{(\ell+s)(\ell-s+1)}}.$ (3b)
The following Lemma (which follows straightforwardly from Definition 2) will be central to the derivation of our addition theorems.
Lemma 4.
$\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)$ $=-\frac{1}{2}\left[{}_{s}\eth\,+{}_{s}{\overline{\eth}}\,\right]{}_{s}Y_{\ell}^{m}(\theta,\phi),$ (4a)
$\frac{\partial}{\partial\phi}{}_{s}Y_{\ell}^{m}(\theta,\phi)$ $=\frac{{\rm{i}}}{2}\sin\theta\left[{}_{s}\eth\,-{}_{s}{\overline{\eth}}\,\right]{}_{s}Y_{\ell}^{m}(\theta,\phi)$
$\qquad-{\rm{i}}s\cos\theta{}_{s}Y_{\ell}^{m}(\theta,\phi).$ (4b)
The well-known usual addition theorem [5, 12, 15, 16] for SWSH depends on the Euler angles $\alpha$, $\beta$, $\gamma$, defined below.
Definition 5.
The Euler angles $\alpha$, $\beta$, $\gamma$ are given in terms of angles $(\theta,\phi)$ and $(\theta^{\prime},\phi^{\prime})$ as follows [15]:
$\cot\alpha$ $=\cos\theta\cot\left(\phi-\phi^{\prime}\right)-\cot\theta^{\prime}\sin\theta\csc\left(\phi-\phi^{\prime}\right),$ (5a)
$\cos\beta$ $=\cos\theta\cos\theta^{\prime}+\sin\theta\sin\theta^{\prime}\cos\left(\phi-\phi^{\prime}\right),$ (5b)
$\cot\gamma$ $=\cos\theta^{\prime}\cot\left(\phi-\phi^{\prime}\right)-\cot\theta\sin\theta^{\prime}\csc\left(\phi-\phi^{\prime}\right).$ (5c)
Remark.
The definitions (5) are, strictly speaking, valid only when $\phi-\phi^{\prime}$ is not a multiple of $\pi$. The corresponding definitions of the Euler angles when $\phi-\phi^{\prime}$ is a multiple of $\pi$ can be found in [15]. For our purposes, we require only that $\alpha=\beta=\gamma=0$ in the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$.
We are now in a position to state the usual addition theorem for SWSH .
Proposition 6.
The addition theorem for spin-weighted spherical harmonics is [15]:
$\left(-1\right)^{s}\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\\ ={\sqrt{\frac{2\ell+1}{4\pi}}}{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}}(\beta,\gamma).$ (6)
Remark.
The addition theorem (6) is valid only when the spins $s$, $s^{\prime}$ differ by an integer, and $\ell\geq\max\{|s|,|s^{\prime}|\}$. We have written (6) in an alternative, but equivalent, form to that displayed in [15]. The proof of Theorem 6 in [15] is presented only for integer spins, but the result is equally valid for half-integer spins and can be straightforwardedly derived using, for example, Definition 1 and the properties of the Wigner $D$-functions [15, 27]. The addition theorem takes a slightly different form in some references [5, 12] due to differences in the definition of the SWSH .
To take the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$ in the addition theorem, we require the coincidence limit of the SWSH :
Lemma 7.
We have [15]
${}_{s}Y_{\ell}^{m}(0,0)$ $=\left(-1\right)^{s}\delta_{s,-m}{\sqrt{\frac{2\ell+1}{4\pi}}},$ (7)
where $\delta_{s,s^{\prime}}$ is the usual Kronecker delta:
$\delta_{s,s^{\prime}}=\begin{cases}1&{\mbox{if $s^{\prime}=s$}},\\ 0&{\mbox{otherwise}}.\end{cases}$ (8)
Corollary 8.
The coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$ of the addition theorem (6) is:
$\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)=\frac{2\ell+1}{4\pi}\delta_{s,s^{\prime}},$ (9)
Our purpose in this note is to derive generalizations of Proposition 6 and Corollary 8 which involve a derivative of either (or both) of the SWSH in the sum.
III Further addition theorems
We now summarize our results, which will be derived in the following section. To the best of our knowledge, these have not previously appeared explicitly in the literature. These results are valid under the same conditions on the quantum numbers $s$, $s^{\prime}$ and $\ell$ as for the original addition theorem, namely the spins $s$, $s^{\prime}$ can only differ by an integer, and it must be the case that $\ell\geq\max\{|s|,|s^{\prime}|\}$.
Proposition 9.
$(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=\frac{1}{2}{\sqrt{\frac{2\ell+1}{4\pi}}}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right\},$ (10a)
$(-1)^{s}\sum_{m=-\ell}^{\ell}m\,{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=-\frac{1}{2}{\sqrt{\frac{2\ell+1}{4\pi}}}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\sin\theta\right.$
$\qquad\qquad\qquad\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\sin\theta+2s\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\cos\theta\right\},$ (10b)
$(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[\frac{\partial}{\partial\theta^{\prime}}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=-\frac{1}{4}{\sqrt{\frac{2\ell+1}{4\pi}}}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad-{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.\qquad-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right],$ (10c)
$(-1)^{s}\sum_{m=-\ell}^{\ell}m^{2}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=-{\sqrt{\frac{2\ell+1}{4\pi}}}\bigg{\{}\frac{1}{4}\sin\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.\qquad+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]$
$\qquad-\frac{s^{\prime}}{2}\sin\theta\cos\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right]$
$\qquad+\frac{s}{2}\cos\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)+{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]$
$\qquad-ss^{\prime}{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\cos\theta\cos\theta^{\prime}\bigg{\}},$ (10d)
$(-1)^{s}\sum_{m=-\ell}^{\ell}m\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=-\frac{1}{4}{\sqrt{\frac{2\ell+1}{4\pi}}}\bigg{\{}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.\qquad+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.\qquad-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]\sin\theta^{\prime}$
$\qquad-2s^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right]\cos\theta^{\prime}\bigg{\}}.$ (10e)
Remark.
The addition theorems (10) all involve one derivative of a SWSH . Similar theorems involving two or more derivatives acting on the same SWSH can be deduced from the above, together with the differential equation governing the SWSH [20, 6, 10, 15, 16]
$\left\{\frac{1}{\sin\theta}\partial_{\theta}\left[\sin\theta\,\partial_{\theta}\right]-\frac{1}{\sin^{2}\theta}\left[s^{2}-2{\rm{i}}s\cos\theta\,\partial_{\phi}-\partial^{2}_{\phi}\right]\right\}{}_{s}Y_{\ell}^{m}(\theta,\phi)=\ell\left[\ell+1\right]{}_{s}Y_{\ell}^{m}(\theta,\phi),$ (11)
and noting that $\partial_{\phi}\left[{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]={\rm{i}}m{}_{s}Y_{\ell}^{m}(\theta,\phi)$.
Taking the coincidence limit $\theta^{\prime}=\theta$, $\phi^{\prime}=\phi$ in each of the results in (10) gives respectively, using (7):
Corollary 10.
$\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)$ $=-\frac{2\ell+1}{8\pi}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,\delta_{s+1,s^{\prime}}-{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,\delta_{s,s^{\prime}+1}\right\},$ (12a)
$\sum_{m=-\ell}^{\ell}m\,{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)$ $=\frac{2\ell+1}{8\pi}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,\delta_{s+1,s^{\prime}}+{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,\delta_{s,s^{\prime}+1}\right\}\sin\theta$
$\qquad-\frac{2\ell+1}{4\pi}s\,\delta_{s,s^{\prime}}\cos\theta,$ (12b)
$\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[\frac{\partial}{\partial\theta}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)\right]$ $=\frac{2\ell+1}{16\pi}\left[2\left(\ell^{2}+\ell-s^{2}\right)\,\delta_{s,s^{\prime}}\right.$
$\left.\qquad-{\sqrt{\left(\ell-s-1\right)\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s+2\right)}}\,\delta_{s+2,s^{\prime}}\right.$
$\left.\qquad-{\sqrt{\left(\ell-s^{\prime}-1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)\left(\ell+s^{\prime}+2\right)}}\,\delta_{s,s^{\prime}+2}\right],$ (12c)
$\sum_{m=-\ell}^{\ell}m^{2}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)$
$=\frac{2\ell+1}{4\pi}\bigg{\{}\left[\frac{1}{2}\left(\ell^{2}+\ell-s^{2}\right)\sin^{2}\theta+s^{2}\cos^{2}\theta\right]\,\delta_{s,s^{\prime}}$
$+\frac{1}{4}\sin^{2}\theta\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,\delta_{s+1,s^{\prime}-1}\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\right]\,\delta_{s-1,s^{\prime}+1}$
$-\frac{1}{2}\sin\theta\cos\theta\left[\left(2s^{\prime}+1\right){\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,\delta_{s,s^{\prime}+1}+\left(2s+1\right){\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,\delta_{s+1,s^{\prime}}\right]\bigg{\}},$ (12d)
$\sum_{m=-\ell}^{\ell}m\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta,\phi)$
$=\frac{2\ell+1}{16\pi}\bigg{\{}\left[-2s\,\delta_{s,s^{\prime}}-{\sqrt{\left(\ell-s^{\prime}-1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)\left(\ell+s^{\prime}+2\right)}}\,\delta_{s,s^{\prime}+2}\right.$
$\left.+{\sqrt{\left(\ell-s-1\right)\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s+2\right)}}\,\delta_{s+2,s^{\prime}}\right]\sin\theta$
$-2s^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,\delta_{s+1,s^{\prime}}-{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,\delta_{s,s^{\prime}+1}\right]\cos\theta\bigg{\}}.$ (12e)
Finally, if we also set the spins equal, $s^{\prime}=s$, the results in Corollary 10 reduce further, giving, respectively:
Corollary 11.
$\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s}Y_{\ell}^{m*}(\theta,\phi)$ $=0,$ (13a)
$\sum_{m=-\ell}^{\ell}m\,\left|{}_{s}Y_{\ell}^{m}(\theta,\phi)\right|^{2}$ $=-\frac{\left(2\ell+1\right)s}{4\pi}\cos\theta,$ (13b)
$\sum_{m=-\ell}^{\ell}\left|\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right|^{2}$ $=\frac{2\ell+1}{8\pi}\left(\ell^{2}+\ell-s^{2}\right),$ (13c)
$\sum_{m=-\ell}^{\ell}m^{2}\,\left|{}_{s}Y_{\ell}^{m}(\theta,\phi)\right|^{2}$ $=\frac{2\ell+1}{8\pi}\left[\left(\ell^{2}+\ell-s^{2}\right)\sin^{2}\theta+2s^{2}\cos^{2}\theta\right],$ (13d)
$\sum_{m=-\ell}^{\ell}m\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s}Y_{\ell}^{m*}(\theta,\phi)$ $=-\frac{\left(2\ell+1\right)s}{8\pi}\sin\theta.$ (13e)
When $s=0$, the results (13) reduce to those found in, for example, App. C of [28] (see also [27, §5.10]).
Remark.
The precise form of the addition theorems (10, 12) depends on the conventions used for the definition of the SWSH . However, the final results in (13) are independent of the choice of phase in the SWSH .
IV Derivation of new addition theorems
Our overall strategy in deriving the results in Proposition 9 is to apply appropriate combinations of the operators ð${}_{s}\eth\,$, ð¯${}_{s}{\overline{\eth}}\,$, ðssubscriptsuperscriptitalic-ð${}_{s^{\prime}}\eth\,^{\prime}$, and ð¯ssubscriptsuperscript¯italic-ð${}_{s^{\prime}}{\overline{\eth}}\,^{\prime}$ (Definition 2) to the original addition theorem (6). We then employ the “raising and lowering” properties of these operators (Proposition 3).
Derivation of (10a).
$(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$ $=-\frac{(-1)^{s}}{2}\sum_{m=-\ell}^{\ell}\left[{}_{s}\eth\,{}_{s}Y_{\ell}^{m}(\theta,\phi)+{}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=-\frac{(-1)^{s}}{2}\sum_{m=-\ell}^{\ell}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{}_{s+1}Y_{\ell}^{m}(\theta,\phi)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=\frac{1}{2}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\sum_{m=-\ell}^{\ell}(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\sum_{m=-\ell}^{\ell}(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right\}$
$=\frac{1}{2}{\sqrt{\frac{2\ell+1}{4\pi}}}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right\}.$ (14)
Derivation of (10b).
$(-1)^{s}\sum_{m=-\ell}^{\ell}m\,{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$ $=-{\rm{i}}\,(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\phi}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=\frac{(-1)^{s}}{2}\sum_{m=-\ell}^{\ell}\left[{}_{s}\eth\,{}_{s}Y_{\ell}^{m}(\theta,\phi)-{}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\sin\theta-s(-1)^{s}\cos\theta\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=\frac{(-1)^{s}}{2}\sum_{m=-\ell}^{\ell}\left\{\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{}_{s+1}Y_{\ell}^{m}(\theta,\phi)+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right\}\sin\theta$
$\qquad-s(-1)^{s}\cos\theta\sum_{m=-\ell}^{\ell}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=\frac{1}{2}\left\{-{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\sum_{m=-\ell}^{\ell}(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\sum_{m=-\ell}^{\ell}(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right\}\sin\theta$
$-s\cos\theta\sum_{m=-\ell}^{\ell}(-1)^{s}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=-\frac{1}{2}{\sqrt{\frac{2\ell+1}{4\pi}}}\left\{{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right\}\sin\theta-s{\sqrt{\frac{2\ell+1}{4\pi}}}\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\cos\theta.$ (15)
Derivation of (10c).
$(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[\frac{\partial}{\partial\theta^{\prime}}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=\frac{(-1)^{s}}{4}\sum_{m=-\ell}^{\ell}\left[{}_{s}\eth\,{}_{s}Y_{\ell}^{m}(\theta,\phi)+{}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[{}_{s^{\prime}}\eth\,^{\prime}{}_{s}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})+{}_{s^{\prime}}{\overline{\eth}}\,^{\prime}{}_{s}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})\right]^{*}$
$=\frac{(-1)^{s}}{4}\sum_{m=-\ell}^{\ell}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{}_{s+1}Y_{\ell}^{m}(\theta,\phi)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]$
×$\times\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})-{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=-\frac{1}{4}\sum_{m=-\ell}^{\ell}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=-\frac{1}{4}{\sqrt{\frac{2\ell+1}{4\pi}}}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.-{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right].$ (16)
Derivation of (10d).
$(-1)^{s}\sum_{m=-\ell}^{\ell}m^{2}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$=(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\phi}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[\frac{\partial}{\partial\phi^{\prime}}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=(-1)^{s}\sum_{m=-\ell}^{\ell}\left\{\frac{{\rm{i}}}{2}\sin\theta\left[{}_{s}\eth\,-{}_{s}{\overline{\eth}}\,\right]{}_{s}Y_{\ell}^{m}(\theta,\phi)-{\rm{i}}s\cos\theta{}_{s}Y_{\ell}^{m}(\theta,\phi)\right\}$
×$\times\left\{\frac{{\rm{i}}}{2}\sin\theta^{\prime}\left[{}_{s^{\prime}}\eth\,^{\prime}-{}_{s^{\prime}}{\overline{\eth}}\,^{\prime}\right]{}_{s^{\prime}}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})-{\rm{i}}s^{\prime}\cos\theta^{\prime}{}_{s^{\prime}}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})\right\}^{*}$
$=(-1)^{s}\sum_{m=-\ell}^{\ell}\bigg{\{}\frac{{\rm{i}}}{2}\sin\theta\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{}_{s+1}Y_{\ell}^{m}(\theta,\phi)+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]$
$-{\rm{i}}s\cos\theta{}_{s}Y_{\ell}^{m}(\theta,\phi)\bigg{\}}$
×$\times\bigg{\{}-\frac{{\rm{i}}}{2}\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})+{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$+{\rm{i}}s^{\prime}\cos\theta^{\prime}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\bigg{\}}$
$=-\sum_{m=-\ell}^{\ell}\bigg{\{}\frac{1}{4}\sin\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$-\frac{s^{\prime}}{2}\sin\theta\cos\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$+\frac{s}{2}\cos\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right](-1)^{s}{}_{s}Y_{\ell}^{m}(\theta,\phi)$
$-ss^{\prime}(-1)^{s}{}_{s}Y_{\ell}^{m}(\theta,\phi){}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\cos\theta\cos\theta^{\prime}\bigg{\}}$
$=-{\sqrt{\frac{2\ell+1}{4\pi}}}\bigg{\{}\frac{1}{4}\sin\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]$
$-\frac{s^{\prime}}{2}\sin\theta\cos\theta^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right]$
$+\frac{s}{2}\cos\theta\sin\theta^{\prime}\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)+{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]$
$-ss^{\prime}{\rm{e}}^{-{\rm{i}}s\alpha}{}_{s}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\cos\theta\cos\theta^{\prime}\bigg{\}}.$ (17)
Derivation of (10e).
$(-1)^{s}\sum_{m=-\ell}^{\ell}m\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})$
$={\rm{i}}\,(-1)^{s}\sum_{m=-\ell}^{\ell}\left[\frac{\partial}{\partial\theta}{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left[\frac{\partial}{\partial\phi^{\prime}}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]$
$=-\frac{{\rm{i}}}{2}(-1)^{s}\sum_{m=-\ell}^{\ell}\left[{}_{s}\eth\,{}_{s}Y_{\ell}^{m}(\theta,\phi)+{}_{s}{\overline{\eth}}\,{}_{s}Y_{\ell}^{m}(\theta,\phi)\right]\left\{\frac{{\rm{i}}}{2}\sin\theta^{\prime}\left[{}_{s^{\prime}}\eth\,^{\prime}-{}_{s^{\prime}}{\overline{\eth}}\,^{\prime}\right]{}_{s^{\prime}}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})-{\rm{i}}s^{\prime}\cos\theta^{\prime}{}_{s^{\prime}}Y_{\ell}^{m}(\theta^{\prime},\phi^{\prime})\right\}^{*}$
$=\frac{(-1)^{s}}{4}\sum_{m=-\ell}^{\ell}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{}_{s+1}Y_{\ell}^{m}(\theta,\phi)-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]$
×$\times\bigg{\{}\left[{\sqrt{\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})+{\sqrt{\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]\sin\theta^{\prime}$
$-2s^{\prime}\cos\theta^{\prime}{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\bigg{\}}$
$=-\frac{1}{4}\sum_{m=-\ell}^{\ell}\bigg{\{}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi)\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\,{}_{s^{\prime}+1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi)\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\,{}_{s^{\prime}-1}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\right]\sin\theta^{\prime}$
$-2s^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,(-1)^{s+1}{}_{s+1}Y_{\ell}^{m}(\theta,\phi)\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,(-1)^{s-1}{}_{s-1}Y_{\ell}^{m}(\theta,\phi)\right]{}_{s^{\prime}}Y_{\ell}^{m*}(\theta^{\prime},\phi^{\prime})\cos\theta^{\prime}\bigg{\}}$
$=-\frac{1}{4}{\sqrt{\frac{2\ell+1}{4\pi}}}\bigg{\{}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell-s^{\prime}\right)\left(\ell+s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}-1}(\beta,\gamma)\right.$
$\left.+{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)\left(\ell+s^{\prime}\right)\left(\ell-s^{\prime}+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s-1}Y_{\ell}^{-s^{\prime}+1}(\beta,\gamma)\right]\sin\theta^{\prime}$
$-2s^{\prime}\left[{\sqrt{\left(\ell-s\right)\left(\ell+s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s+1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right.$
$\left.-{\sqrt{\left(\ell+s\right)\left(\ell-s+1\right)}}\,{\rm{e}}^{-{\rm{i}}(s-1)\alpha}{}_{s+1}Y_{\ell}^{-s^{\prime}}(\beta,\gamma)\right]\cos\theta^{\prime}\bigg{\}}.$ (18)
V Conclusions
In this paper we have presented some new addition theorems for SWSH , generalizing the well-known addition theorem (6) [5, 12, 15, 16]. These new results, like the original addition theorem, involve a sum over the azimuthal quantum number $m$, and the summand involves two SWSH , with a single derivative acting on one (or both) of these. Since the SWSH satisfy a second order differential equation (11), similar addition theorems involving two or more derivatives acting on a SWSH can easily be derived. The derivation of our new results follows from a straightforward application of the raising and lowering operators ð${}_{s}\eth\,$ and ð¯${}_{s}{\overline{\eth}}\,$ (3) to the original addition theorem.
Our original motivation for deriving these results was to find the sums (13) which we required for applications in quantum field theory on black hole space-times [29, 30]. However, our main results (10) are much more general and we hope that by presenting them here they will be useful to researchers in wider application areas.
Acknowledgements.
The work of A.M. is supported by an EPSRC studentship. The work of E.W. is supported by STFC grant number ST/X000621/1. No data were created or analyzed in this study.